2:x = X:50 what one number can replace x
Answer by jim_thompson5910(35256)

2:x = x:50

2/x = x/50

2*50 = x*x

100 = x^2

x^2 = 100

x = +-sqrt(100)

x = sqrt(100) or x = -sqrt(100)

x = 10 or x = -10

So x could be 10 or -10